# An Overview of CV-MDI-QKD

### v1

@@ Abstract
As quantum key distribution (QKD) emerges as a robust defense against quantum computer threats, significant advancements have been realized by researchers. A pivotal focus has been the development of protocols that not only simplify hardware implementation like the use of continuous-variable (CV) systems, but also negate the necessity for trusted nodes, as seen with the measurement-device independent (MDI) approach. This paper delves into the integration of these methodologies in the CV-MDI-QKD protocol, offering an in-depth exploration of its evolution, primary characteristics, and the latest advancements in both theory and experiment.
@@ I Introduction
To address (i), one such milestone was the derivation of the Pirandola-Laurenza-Ottaviani-Banchi (PLOB) bound Pirandola et al. (2017) which establishes the exact fundamental trade-off between communication rate and distance. This limit puts an upper bound on all point-to-point implementations of QKD. Namely, for a lossy bosonic channel, which could underlie models of optical-fibre and free-space channels, two users cannot distribute more than the secret key capacity of the link $-\log_{2}(1-\tau)$, where $\tau$ quantifies channel’s loss. The PLOB bound provides a guide for all point-to-point QKD protocols, outlining an optimal performance level that can guide future protocol/experimental designs.
@@
Nevertheless, the development of repeater-assisted protocols invites further questions. In Refs. Pirandola (2019a, b), the end-to-end capacities are shown to be achievable by assuming either trusted QKD repeaters or entanglement-based repeaters. In a practical QKD network, it would be ideal to use repeaters simultaneously untrusted, so they could still be usable even if compromised by an attacker, and not based on entanglement distribution, so they can be more easily deployed in realistic technological environments. The first step done in this direction was the introduction of measurement device independent (MDI) QKD Braunstein and Pirandola (2012); Lo et al. (2012); Liu et al. (2013); Curty et al. (2014) showing that QKD is still possible in a single-hop chain where Alice and Bob connects to an untrusted, and entanglement-free, repeater. More recently, the MDI idea was extended to design the twin-field QKD protocol Lucamarini et al. (2018); Minder et al. (2019), which can also exceed the PLOB bound.
Soon after the introduction of the MDI protocol, a CV version was designed and and experimentally realized Pirandola et al. (2013, 2015). CV-MDI-QKD aims to achieve the same features as DV-MDI-QKD while ensuring much higher key rates at short distances. Many works followed proposing variants and improvements. Several works implemented techniques based on photon subtraction Zhao et al. (2018); Ma et al. (2018), noiseless linear amplifiers Zhang et al. (2015) and phase-sensitive amplifiers Wang et al. (2019). Attempts have been made on the encoding side by replacing coherent states with squeezed or thermal states Zhang et al. (2014); Bai et al. (2019a, b), and replacing Gaussian modulation with discrete modulation or dual-phase modulation Ma et al. (2019); Liao et al. (2018).
@@
Figure 1: Schematic of CV-MDI-QKD. In the entanglement-based description, Alice and Bob transmit one mode ($A$ and $B$) of their TMSV states, denoted by red circles. In the equivalent and practical prepare-and-measure description (dashed boxes), Alice and Bob directly send Gaussian-modulated coherent states to the relay. Eve has a pair of TMSV states, denoted by grey circles. Her modes $E_{1}$ and $E_{2}$ are made to interact with Alice and Bob’s signal modes via beam splitters of transmissivities $\tau_{A}$ and $\tau_{B}$, respectively (each of these interactions corresponds to an ‘entangling cloner’). Eve’s output modes after the beamsplitters and her retained modes ($e_{1}$ and $e_{2}$) are stored in a quantum memory for collective measurement at the end of the protocol (this is known as a ‘collective attack’). Alice’s and Bob’s output modes $A^{\prime}$ and $B^{\prime}$ are mixed in a $1/2$ beam splitter at the relay, whose outputs are subsequently measured using two conjugate homodyne detections with outcomes $q_{-}$ and $p_{+}$, respectively. These real outcomes form a single complex value $\gamma$ that is publicly broadcast.
This review is structured as follows. In Sec. II we provide a schematic overview of the protocol. In Sec. III, we discuss the secret key rate of CV-MDI-QKD based on Gaussian modulation and its post-selection variant. In Sec. IV, we review finite-size and composability. In Sec. VI we discuss the maib experimental implementations of CV-MDI-QKD. In Sec. V, we weigh up quantum network complexity, including CV-MDI-QKD star and three-node networks. We summarize our manuscript in Sec. VII.
@@ II Features of the protocol
b. Alice and Bob send their Gaussian-modulated coherent states $|\alpha\rangle$ and $|\beta\rangle$ to the relay through quantum channels that could be, e.g., optical fibres or free-space links. These channels are theoretically simulated via beam splitters with transmissivities $\tau_{A}$ and $\tau_{B}$. The eavesdropper (Eve) can attack the links by overlapping one mode of her TMSV states at the beam splitters. See the next section for more details.
c. At the relay, Charlie uses a CV Bell measurement to perform a joint detection. This step includes overlapping signals received from Alice and Bob’s station on a balanced beam splitter and homodyning both outlet beams. Precisely, Charlie measures orthogonal quadratures $q$ and $p$ obtaining $q_{-}$ and $p_{+}$, respectively. The resultant complex value $\gamma=q_{-}+ip_{+}$ is then broadcast to Alice and Bob.
d. Alice and Bob post-process their measured/modulation data along with the corresponding relay outcomes. The post-processing includes channel parameter estimation (PE), error correction (EC), and privacy amplification (PA). These operations will provide the final secret shared key.
@@ III Security analysis
The two end users, Alice and Bob, each have access to independent, identical zero-mean bi-variate Gaussian distributions with the same variance $\mu-1$ (assumed to be large). From these distributions, they draw two amplitude values $\alpha$ and $\beta$ therefore preparing Gaussian-modulated coherent states |$\ket{\alpha}$ and |$\ket{\beta}$ respectively (see prepare-and-measure description in Fig. 1). These two states are sent to a central detection station, which is in principle controlled by Eve. Operated correctly, the station performs a Bell detection in which the states are combined on a balanced beamsplitter and undergo conjugate homodyne detections. The results of these measurements are combined into a single complex variable $\gamma$ which is publicly broadcast. It is clear that initially the mutual information $I(\alpha:\beta)=0$ since the coherent states are prepared from independent Gaussian distributions. However after the announcement of $\gamma$, Alice and Bob are able to decode each other’s variable and thus $I(\alpha:\beta|\gamma)>0$. Despite having access to $\gamma$, Eve is unable to infer the parties’ variables, i.e., $I(E:\alpha)=I(E:\beta)=0$. Eve is therefore forced to attack the channels or alter the operation of the relay, both of which may be detected by Alice and Bob in post-processing.
@@
where $\bm{Z}=\text{diag}(1,-1)$. Mode $a$ undergoes heterodyne detection, projecting the travelling mode $A$ into the coherent state |$\ket{\alpha}$, the modulation of which is in one-to-one correspondence with the measurement outcome $\tilde{\alpha}$. In the limit of large modulation $\mu\rightarrow\infty$, we have $\alpha=\tilde{\alpha}^{*}$. The one-to-one correspondence between $\alpha$ and $\tilde{\alpha}$ renders the two variables informationally equivalent. Symmetric description holds for Bob. Alice’s and Bob’s modes $A$ and $B$ are sent to the central relay for Bell detection and, by the commutation of local measurements, Alice’s and Bob’s heterodyne detections of their retained modes $a$ and $b$ may be delayed until the end of the protocol. The protocol can thus be seen as an entanglement swapping protocol, generating an entangled Gaussian state between Alice’s and Bob’s retained modes by the Bell detection of the transmitted modes. The entanglement-based representation is a useful mathematical tool for studying the security of the protocol but also provides a potential version of it that can prevent side-channel attacks on state preparation Braunstein and Pirandola (2012).
@@ III.1 Gaussian attacks and asymptotic key rate
In the symmetric configuration, Alice and Bob are equidistant from the relay. In this case, the rate-distance scaling is shown in Fig. 2. The distance is calculated by assuming the parties connected to the relay by fibre optic links with a loss of $0.2$ dB/km. That is the transmissivity of the beamsplitters used in Eve’s attack may be given by $\tau_{i}=10^{-0.02d}$ where $d$ is the distance in kilometers of the link.
Figure 2: Asymptotic key rate (bits/use) of the symmetric CV-MDI-QKD protocol versus fibre distance ($0.2$dB/km) for various values of thermal noise $\bar{n}$ affecting each link. The solid black line indicates the rate under pure loss. All rates are optimized with large modulation. Unit reconciliation and perfect detection efficiency are assumed. Figure adapted from Ref. Fletcher (2023)
@@
In Fig. 3 we show the rate-distance scaling in the asymmetric configuration, where we assume that Alice is the encoder and Bob the decoder (this means that Bob needs to guess Alice’s variable). If we fix the distance of Alice at $5$m from the relay, we can see that Bob can be very far. The distance scaling is much improved in this configuration, enabling communication at over $100$km between the two end parties. Indeed, it is shown in Ref. Pirandola et al. (2015), that if Bob’s channel is affected by pure loss only, in the limit $\tau_{A}\rightarrow 1$ (Alice brought arbitrarily close to the relay) that the rate only goes to zero for $\tau_{B}\rightarrow 0$, which represents Bob arbitrarily far away.
The reverse scenario in which Bob’s position is fixed close to the relay is substantially less effective. Even with $\tau_{B}\rightarrow 1$ and pure loss in Alice’s channel, a positive rate can only be maintained to $6.8$km. This difference of performance is related to the distinction between direct and reverse reconciliation in CV QKD. In the limit $\tau_{A}\rightarrow 1$ in which Alice (the encoder) approaches the relay, we recover a point-to-point no-switching protocol in reverse reconciliation. Conversely, with $\tau_{B}\rightarrow 1$ (decoder close to the relay), we have a point-to-point no-switching protocol in direct reconciliation.
Figure 3: Asymptotic key rate (bits/use) of CV-MDI-QKD in the asymmetric configuration versus fibre distance ($0.2$dB/km) between Alice and Bob. Black lines refer to Alice at a fixed distance of $5$m from the relay, while Bob is at a variable fibre distance. The red lines represent the opposite scenario in which Bob’s is fixed at $5$m from the relay and Alice’s fibre distance is allowed to vary. Solid lines indicate the rates under pure loss and the dashed with thermal noise equal to $\bar{n}=0.05$. All rates are optimized with large modulation and perfect detection efficiency is assumed. Figure adapted from Ref. Fletcher (2023)
@@ III.2 Post-selection
Ref. Wilkinson et al. (2020) introduced post-selection for CV-MDI-QKD. In the prepare-and-measure formalism, Alice and Bob again prepare coherent states and each has access to zero-mean bi-variate Gaussian distributions, from which they draw two real variables for the values of the displacement of the $q$ and $p$ quadratures of the coherent states they will prepare. They encode the magnitude of their value for the $q$ displacement as $\mathbb{Q}_{A}$ and $\mathbb{Q}_{B}$ and the sign of the value as $\kappa_{A}$ and $\kappa_{B}$. Similarly, the absolute values drawn for the $p$ displacements are encoded as $\mathbb{P}_{A}$ and $\mathbb{P}_{B}$, and the signs as $\kappa_{A}^{\prime}$ and $\kappa_{B}^{\prime}$. Thus, Alice and Bob prepare the coherent states |$\ket{\alpha=\frac{1}{2}(\kappa_{A}\mathbb{Q}_{A}+i\kappa_{A}^{\prime}\mathbb{P}_{A})}$ and |$\ket{\beta=\frac{1}{2}(\kappa_{B}\mathbb{Q}_{B}+i\kappa_{B}^{\prime}\mathbb{P}_{B})}$, respectively.
@@
Under phase-insensitive collective Gaussian attacks (such as the typical case of a collective attack based on two entangling cloners), the two quadratures are symmetric in terms of the rate so, for simplicity, we shall consider only the use of the $q$ quadrature. The measurement yields the distribution $p(\gamma_{q}|\kappa_{A},\kappa_{B},\mathbb{Q}_{A},\mathbb{Q}_{B})$ (note we have implicitly removed conditioning from variables on which there is no dependence) from which Bayes’ theorem may be used to calculate $p(\kappa_{A}|\kappa_{B},\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and $p(\kappa_{A},\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and hence determine the single point mutual information $\tilde{I}(\kappa_{A}:\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$. Similarly we can the single point Holevo information $\tilde{\chi}(E:\kappa_{A}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and the single-point rate
$\tilde{R}(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ $=\tilde{I}(\kappa_{A}:\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$
$-\tilde{\chi}(E:\kappa_{A}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q}),$ (3)
@@
[…]
@@
[…]
Figure 4: Comparative rate-distance performance of CV-MDI-QKD with and without post-selection in the symmetric configuration. The black lines indicate the original CV-MDI-QKD protocol Pirandola et al. (2015) and the red the equivalent protocol with post-selection Wilkinson et al. (2020). Solid lines are the rates under pure loss and dashed lines refer to thermal noise equal to $\bar{n}=0.05$ thermal photons. All the rates (bits/use) are optimized over the modulation of the prepared coherent states and assume perfect reconciliation and detection efficiency. Figure adapted from Ref. Fletcher (2023)

### v3

# thanks: Corresponding author

@@ Abstract
As quantum key distribution (QKD) emerges as a robust defense against quantum computer threats, significant advancements have been realized by researchers. A pivotal focus has been the development of protocols that not only simplify hardware implementation, such as the use of continuous-variable (CV) systems, but also eliminate the necessity for trusted nodes, as seen with the measurement-device-independent (MDI) approach. This paper delves into the integration of these methodologies in the CV-MDI-QKD protocol, offering an in-depth exploration of its evolution, primary characteristics, and the latest advancements in both theory and experiment.
@@ I Introduction
To address (i), one such milestone was the derivation of the Pirandola-Laurenza-Ottaviani-Banchi (PLOB) bound Pirandola et al. (2017) which establishes the exact fundamental trade-off between communication rate and distance. This limit puts an upper bound on all point-to-point implementations of QKD. Namely, for a lossy bosonic channel, which could underlie models of optical-fibre and free-space channels, two users cannot distribute more than the secret key capacity of the link $\log_{2}(1-\tau)$, where $\tau$ quantifies channel’s loss. The PLOB bound provides a guide for all point-to-point QKD protocols, outlining an optimal performance level that can guide future protocol/experimental designs.
@@
Nevertheless, the development of repeater-assisted protocols invites further questions. In Refs. Pirandola (2019a, b), the end-to-end capacities are shown to be achievable by assuming either trusted QKD repeaters or entanglement-based repeaters. In a practical QKD network, it would be ideal to use repeaters simultaneously untrusted, so they could still be usable even if compromised by an attacker, and not based on entanglement distribution, so they can be more easily deployed in realistic technological environments. To address (ii) and (iii) simultaneously, the first step done in this direction was the introduction of measurement device independent (MDI) QKD Braunstein and Pirandola (2012); Lo et al. (2012); Liu et al. (2013); Curty et al. (2014) showing that QKD is still possible in a single-hop chain where Alice and Bob connects to an untrusted, and entanglement-free, repeater. More recently, the MDI idea was extended to design the twin-field QKD protocol Lucamarini et al. (2018); Minder et al. (2019), which can also exceed the PLOB bound.
Soon after the introduction of the MDI protocol, a CV version was designed and experimentally realized Pirandola et al. (2013, 2015). CV-MDI-QKD aims to achieve the same features as DV-MDI-QKD while ensuring much higher key rates at short distances. Many works followed proposing other studies, variants and improvements. Following Ref. Pirandola et al. (2013), CV-MDI-QKD was also investigated in Refs. Ma et al. (2014); Li et al. (2014) with some inaccuracies. Several works implemented techniques based on photon subtraction Zhao et al. (2018); Ma et al. (2018), noiseless linear amplifiers Zhang et al. (2015), phase-sensitive amplifiers Wang et al. (2019), and in free-space settings Ghalaii and Pirandola (2023). Attempts have been made on the encoding side by replacing coherent states with squeezed or thermal states Zhang et al. (2014); Bai et al. (2019a, b), and replacing Gaussian modulation with discrete modulation or dual-phase modulation Ma et al. (2019); Liao et al. (2018).
@@
Figure 1: Schematic of CV-MDI-QKD. In the entanglement-based description, Alice and Bob transmit one mode (AA and $B$) of their TMSV states, denoted by red circles. In the equivalent and practical prepare-and-measure description (dashed boxes), Alice and Bob directly send Gaussian-modulated coherent states to the relay. Eve has a pair of TMSV states, denoted by grey circles. Her modes $E_{1}$ and $E_{2}$ are made to interact with Alice and Bob’s signal modes via beam splitters of transmissivities $\tau_{A}$ and $\tau_{B}$, respectively (each of these interactions corresponds to an ‘entangling cloner’). Eve’s output modes after the beamsplitters and her retained modes ($e_{1}$ and $e_{2}$) are stored in a quantum memory (QM) for collective measurement at the end of the protocol (this is known as a ‘collective attack’). Alice’s and Bob’s output modes $A^{\prime}$ and $B^{\prime}$ are mixed in a $1/2$ beam splitter at the relay, whose outputs are subsequently measured using two conjugate homodyne detections with outcomes $q_{-}$ and $p_{+}$, respectively. These real outcomes form a single complex value $\gamma$ that is publicly broadcast.
This review is structured as follows. In Sec. II we provide a schematic overview of the protocol. In Sec. III, we discuss the secret key rate of CV-MDI-QKD based on Gaussian modulation and its post-selection variant. In Sec. IV, we review finite-size and composability. In Sec. V, we weigh up quantum network complexity, including CV-MDI-QKD star and three-node networks. In Sec. VI we discuss the main experimental implementations of CV-MDI-QKD. We summarize our manuscript in Sec. VII.
@@ II Features of the protocol
b. Alice and Bob send their Gaussian-modulated coherent states $|\alpha\rangle$ and $|\beta\rangle$ to the relay through quantum channels that could be, e.g., optical fibres or free-space links. These channels can be simulated via beam splitters with transmissivities $\tau_{A}$ and $\tau_{B}$. The eavesdropper (Eve) can attack the links by overlapping one mode of her TMSV states at the beam splitters. This ‘entangling-cloner attack’ represents the most relevant Gaussian attack. See the next section for more details.
c. At the relay, Charlie uses a CV Bell measurement to perform a joint detection. This step includes overlapping signals received from Alice and Bob’s station on a balanced beam splitter and homodyning both outlet beams. Precisely, Charlie measures orthogonal quadratures $q$ and $p$ obtaining $q_{-}$ and $p_{+}$, respectively. The resultant complex value $\gamma=q_{-}+ip_{+}$ is then broadcast to Alice and Bob. The outcome $\gamma$ establishes a-posteriori correlations between the variables of the parties. Consequently, knowing $\gamma$ allows each party to infer the variable of the other party through post-processing.
d. Alice and Bob post-process their measured/modulation data along with the corresponding relay outcomes, with the effect of correlating their data. Next, the post-processing includes channel parameter estimation (PE), error correction (EC), and privacy amplification (PA). These operations will provide the final secret shared key.
@@ III Security analysis
[…]
@@
[…]
@@ III.1 Gaussian attacks and asymptotic key rate
In the symmetric configuration, Alice and Bob are equidistant from the relay. In this case, the rate-distance scaling is shown in Fig. 2. The distance is calculated by assuming the parties connected to the relay by fibre optic links with a loss of $0.2$ dB/km. That is the transmissivity of the beamsplitters used in Eve’s attack may be given by $\tau_{i}=10^{-0.02d_{i}}$ where $d_{i}$ is the distance in kilometers of the links to Eve.
Figure 2: Asymptotic key rate (bits/use) of the symmetric CV-MDI-QKD protocol versus fibre distance (0.20.2dB/km) for various values of thermal noise $\bar{n}$ affecting each link. The solid black line indicates the rate under pure loss. All rates are optimized with large modulation. Unit reconciliation and perfect detection efficiency are assumed. Figure adapted from Ref. Fletcher (2023).
@@
In Fig. 3 we show the rate-distance scaling in the asymmetric configuration, where we assume that Alice is the encoder and Bob the decoder (this means that Bob needs to guess Alice’s variable). If we fix the distance of Alice at 55m from the relay, we can see that Bob can be very far. The distance scaling is much improved in this configuration, enabling communication at over 100100km between the two end parties. Indeed, it is shown in Ref. Pirandola et al. (2015), that if Bob’s channel is affected by pure loss only, in the limit $\tau_{A}\rightarrow 1$ (Alice brought arbitrarily close to the relay) that the rate only goes to zero for $\tau_{B}\rightarrow 0$, which represents Bob arbitrarily far away.
The reverse scenario in which Bob’s position is fixed close to the relay is substantially less effective. Even with $\tau_{B}\rightarrow 1$ and pure loss in Alice’s channel, a positive rate can only be maintained to 6.86.8km. This difference of performance is related to the distinction between direct and reverse reconciliation in CV QKD. In the limit $\tau_{A}\rightarrow 1$ in which Alice (the encoder) approaches the relay, we recover a point-to-point no-switching protocol in reverse reconciliation. Conversely, with $\tau_{B}\rightarrow 1$ (decoder close to the relay), we have a point-to-point no-switching protocol in direct reconciliation.
Figure 3: Asymptotic key rate (bits/use) of CV-MDI-QKD in the asymmetric configuration versus fibre distance (0.20.2dB/km) between Alice and Bob. Black lines refer to Alice at a fixed distance of 55m from the relay, while Bob is at a variable fibre distance. The red lines represent the opposite scenario in which Bob’s is fixed at 55m from the relay and Alice’s fibre distance is allowed to vary. Solid lines indicate the rates under pure loss and the dashed with thermal noise equal to $\bar{n}=0.05$. All rates are optimized with large modulation and perfect detection efficiency is assumed. Figure adapted from Ref. Fletcher (2023).
@@ III.2 Post-selection
[…]
@@
Under phase-insensitive collective Gaussian attacks (such as the typical case of a collective attack based on two entangling cloners), the two quadratures are perturbed symmetrically and remain uncorrelated. For this reason, we shall consider only the use of the $q$ quadrature. The measurement yields the distribution $p(\gamma_{q}|\kappa_{A},\kappa_{B},\mathbb{Q}_{A},\mathbb{Q}_{B})$ (note we have implicitly removed conditioning from variables on which there is no dependence) from which Bayes’ theorem may be used to calculate $p(\kappa_{A}|\kappa_{B},\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and $p(\kappa_{A},\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and hence determine the single point mutual information $\widetilde{I}(\kappa_{A}:\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$. Similarly we can the single point Holevo information $\widetilde{\chi}(E:\kappa_{A}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ and the single-point rate
$\widetilde{R}(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$ $=\widetilde{I}(\kappa_{A}:\kappa_{B}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})$
$-\widetilde{\chi}(E:\kappa_{A}|\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q}),$ (3)
@@
$\int\mathrm{max}\big{[}\widetilde{R}(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q}),0\big{]}\leavevmode\nobreak\ p(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})\leavevmode\nobreak\ d\mathbb{Q}\leavevmode\nobreak\ d\gamma_{q}.$ (4)
@@
$R=\int_{\Gamma}p(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma_{q})\leavevmode\nobreak\ \widetilde{R}(\mathbb{Q}_{A},\mathbb{Q}_{B},\gamma)\leavevmode\nobreak\ d\mathbb{Q}\leavevmode\nobreak\ d\gamma_{q}.$ (5)
Figure 4: Comparative rate-distance performance of CV-MDI-QKD with and without post-selection in the symmetric configuration. The black lines indicate the original CV-MDI-QKD protocol Pirandola et al. (2015) and the red the equivalent protocol with post-selection Wilkinson et al. (2020). Solid lines are the rates under pure loss and dashed lines refer to thermal noise equal to $\bar{n}=0.05$ thermal photons. All the rates (bits/use) are optimized over the modulation of the prepared coherent states and assume perfect reconciliation and detection efficiency. Figure adapted from Ref. Fletcher (2023).